OpenStudy (anonymous):

4(3z-1)-3(z+5)=6(z+1)

OpenStudy (anonymous):

Distributive property

OpenStudy (anonymous):

4(3z-1)-3(z+5)=6(z+1) would become 12z-4-3z-15=6z+6 Now combine like terms: 9z-19=6z+6 Edit equation so each group of like terms are in separate sides: 9z=6z+25.....because 19 is added to 6 Then.... 3z=25......because 6z is subtracted from 9z Finally... z=25/3 or 8.33....because 3 is divided by 25 to find x. Hope it helped:)